Circular No. 8518

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

VARIABLE OBJECT IN BOOTES Further to IAUC 8515, T. Puckett and E. Briggs report the discovery

of a variable object (mag 16.0) on an unfiltered CCD frame taken with the 0.50-m automated supernova patrol telescope on Apr. 20.25 and 21.21 UT. The new object is located at α = 14^{h}54^{m}20^{s}.91, δ = +16^{o}24′23′′.8 (equinox 2000.0), which is 37′′ west and 184′′ north of the center of IC 4516, and it was also visible at mag 17.6 on Puckett’s images from Apr. 5. Nothing is visible at this location on images taken by Puckett on 2002 Feb. 13 and 2003 Mar. 15 (limiting mag ∼ 20.0), but a faint, starlike object is visible at this location on Palomar Sky Survey plates taken on 1993 Mar. 23 (blue mag 18.6) and 1991 May 16 (red mag 19.4). H. Yamaoka, Kyushu University, reports a confirmation of this object on an unfiltered CCD image taken by K. Itagaki (Teppo-cho, Yamagata, Japan; 0.60-m telescope) at red mag 15.9 on Apr. 21.577, providing position end figures 20^{s}.85, 24′′.4. Yamaoka adds that this appears to be 3C 306 (possibly a quasar), with the location being very close to a dumbbell-shaped radio source in the NVSS and FIRST catalogues; it also appears to be listed in the USNO-B1.0 catalogue (with position end figures 20^{s}.85, 24′′.7) as having red mag 19.9.

SUPERNOVA 2005bt IN UGC 8205 Further to IAUC 8476, O. Trondal, P. Holvorcem, and M. Schwartz

report their discovery of an apparent supernova (mag ∼ 18.0) on unfiltered Tenagra II 0.81-m telescope images taken on Apr. 20.40 and 21.27 UT. SN 2005bt is located at α = 13^{h}07^{m}13^{s}.07, δ = +58^{o}08′08′′.4 (equinox 2000.0), which is 4′′.4 west and 2′′.2 north of the nucleus of UGC 8205. Nothing was visible at this position in earlier images taken by T. Boles on 2003 May 25.96 (0.35-m reflector; limiting mag ∼ 19.0) and W. Li on 2002 June 19.22 (KAIT; limiting mag ∼ 19.0). T. Puckett provides the following unfiltered CCD magnitudes for SN 2005bt: 2002 Apr. 18, [20; 2003 Apr. 12, [20; 2005 Mar. 19, 18.4; Apr. 17, 18.0; 21, 18.0.

COMET C/2003 T4 (LINEAR) Total visual magnitude estimates: Mar. 10.22 UT, 8.4 (J. J. Gonzalez,

León, Spain, 7×50 binoculars); 17.14, 7.9 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); Apr. 2.86, 8.0 (A. Pearce, Noble Falls, W. Australia, 20×80 binoculars); 17.77, 8.1 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars).

2005 April 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green